OpenStudy (anonymous):

A can in the shape of a right circular cylinder is required to have a volume of 700 cubic centimeters. The top and bottom are made up of a material that costs 6¢ per square centimeter, while the sides are made of material that costs 5¢ per square centimeter. Find a function that describes the total cost of the material as a function of the radius r of the cylinder.

OpenStudy (anonymous):

\[volume of cylinder=\pi r ^{2}h,\] find h \[curved area=2 \pi r *h and then cost\] \[Find area of \top and bottom= 2*\pi r ^{2} and then cost add the two and get the answer.\]

OpenStudy (jdoe0001):

 so the 5cents Area, is really the sides of the can and the 6cents Area, is really just the "circles" atop and bottom \(\bf \textit{lateral area} = 2\pi \times radius \times height\\ \textit{circle's area} = \pi\times radius^2\\ f(x) = \textit{lateral area + circle's area + circle's area}\)